Suppose thatr an athlete free dives to an elevation of -42 meters in 60 seconds. Find the average rate of change in the divers elevation?
Answer by stanbon(48545)

Suppose thatr an athlete free dives to an elevation of -42 meters in 60 seconds. Find the average rate of change in the divers elevation?
-----------------
Average rate = -42/60 = -7/10 meters per second
=====================================================
Cheers,
Stan H.